rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#8760 — FS#12627 — Bordeaux <> Marseille

Attached to Project— Network
Maintenance
Backbone EU
CLOSED
100%
There's 8x10 down between Bordeaux Marseille from 16/02 to 6:00 GMT, on 2 different providers (incident on the optical transmission).

Update 3 (UTC/CET(CEST)): 16/02/2015 08:38/ 09:38

Engineers are already on site. Incident is localized on the highway. Due to nature of the incident, access to exact break location is restricted by local authorities.
Date:  Wednesday, 18 February 2015, 10:22AM
Reason for closing:  Done
Comment by OVH - Monday, 16 February 2015, 10:33AM

There's little traffic on the POP, and the traffic is rerouting via Paris directly via other secure connections.


Comment by OVH - Monday, 16 February 2015, 15:17PM

The fault is localised in Toulouse. The optic cables melted following an accident/fire.

Preparation of work is in progress. We hope to get services back up before this evening.


Comment by OVH - Monday, 16 February 2015, 15:18PM

The fault has been further explained as vandalism. It appears that a canal was intentionally set on fire causing damages to two separate cables. The fiber provider's field crew is working to run three new 100 meter long cables and install new splice enclosure to begin the required repairs. The fiber provider anticipates fiber splicing to commence at approximately 18:00 GMT.


Comment by OVH - Monday, 16 February 2015, 16:31PM

Permission to access the break location has been granted by the local authorities and preparatory works are currently ongoing.


Comment by OVH - Tuesday, 17 February 2015, 09:39AM

For security reasons, restoration works have stopped overnight and will not resume untill 0700 UTC. No police protection possible.


Comment by OVH - Tuesday, 17 February 2015, 09:51AM

Restoration works are already in progress. We are chasing the engineers for ETR.


Comment by OVH - Tuesday, 17 February 2015, 12:47PM

Half of our connections remounted at 09:43 GMT.

We still need to remount 4x10G.

We're waiting for information from our second provider.


Comment by OVH - Tuesday, 17 February 2015, 14:14PM

Restoration works continue. The estimated time of repair is 19:00 UTC/20:00 CET


Comment by OVH - Wednesday, 18 February 2015, 10:22AM

Network provider has confirmed that the incident is resolved with a permanent fix.